OpenStudy (anonymous):

A total of tickets were sold for the school play. They were either adult tickets or student tickets. There were more student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (texaschic101):

you need to read this question again....there are numbers missing. Please re-post the question

OpenStudy (anonymous):

ya he is right

OpenStudy (anonymous):

A total of 801 tickets were sold for the school play. They were either adult tickets or student tickets. There were 51 more student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (anonymous):

let student ticket x => for adult x-51 => x+x-51=801 => x=750/2 => for adult 750/2-51

OpenStudy (texaschic101):

a + s = 801 s = a + 51 now sub a + 51 in for s in the 1st equation a + a + 51 = 801 2a + 51 = 801 -- subtract 51 from both sides 2a = 801 - 51 2a = 750 -- divide both sides by 2 a = 750/2 a = 375 adult tickets sole were 375